The perimeter of a rectangle with length twice its width is equal to twenty-four. Find the measure of the length and width
Answer by cleomenius(959)

2(2x + x) = 24
4x + 2x = 24
6x = 24
x = 4
4, 8
2*4 + 2*8 = 24 checks.
Cleomenius